OpenStudy (anonymous):

HELP. 2x-2y/3x+3y * 6x+6y/x^2-y^2 Anwsers: 1. 4 2. 2/3(x-y) 3. 4/x+y

OpenStudy (anonymous):

Option 3.

OpenStudy (anonymous):

what about this one? x^2 - 9/ x^2 + 5x + 6 * 1/ X+3 Anwsers: 1. 1 2. 1/x+2 3. x-3/(x+2)(x+3)

OpenStudy (anonymous):

Just a second...

OpenStudy (anonymous):

Option 3.

OpenStudy (campbell_st):

the answer is D \[2(x -y)/3x+3y) \times 2(3x + 3y)/(x-y)(x+y) = ( 2 \times 2)(x+y)\]